# Gradient Interference-Aware Graph Coloring for Multitask Learning

### v1

@@ Abstract
When different objectives conflict with each other in multi-task learning, gradients begin to interfere and slow convergence, thereby reducing the final model’s performance. To address this, we introduce a scheduler that computes gradient interference, constructs an interference graph, and then applies greedy graph-coloring to partition tasks into groups that align well with each other. At each training step, only one group (color class) of tasks are activated. The grouping partition is constantly recomputed as task relationships evolve throughout training. By ensuring that each mini-batch contains only tasks that pull the model in the same direction, our method improves the effectiveness of any underlying multi-task learning optimizer without additional tuning. Since tasks within these groups will update in compatible directions, model performance will be improved rather than impeded. Empirical results on six different datasets show that this interference-aware graph-coloring approach consistently outperforms baselines and state-of-the-art multi-task optimizers.
@@ 1 Introduction
As modern applications increasingly span vision, language, speech, and beyond (generating massive heterogeneous datasets) Kokkinos2016UberNet ; Baltrusaitis2018 ; Liang2022Foundations there is a growing need for models that can learn from and operate across multiple tasks simultaneously. Multi-task learning (MTL) trains a single model to solve several tasks at once, sharing knowledge across them to use data and compute more efficiently Caruana1997 . Early studies showed that this idea can boost accuracy on small vision and language problems, but they also revealed major drawbacks Sener2018 ; Shi2023Recon . When two tasks push the shared network in opposite directions their gradients clash, slowing or even reversing learning. Classic fixes, like manual loss weights or static task schedules Argyriou2007MultiTaskFeature ; AndoZhang2005 ; Evgeniou2005 ; Kang2011 help only in narrow settings and must be tuned for every new dataset Yu2021RevisitingLossWeight .
Recent work has shifted towards gradient-level solutions. Projection methods such as PCGrad Yu2020PCGrad remove the part of one task’s gradient that conflicts with another. Adaptive-rate approaches like AdaTask Yang2023AdaTask shrink the learning rate for any task whose gradients grow too large. These approaches reduce interference in each update, yet they still mix all tasks together every step. As a result, strongly negative pairs continue to impact progress Shi2023Recon ; Li2025PiKE , and the solutions themselves become noisy because task relationships change during training.
Our work tackles this problem by adjusting when tasks are trained rather than how their gradients are modified. We propose a lightweight scheduler that measures recent gradient conflict, builds a conflict graph, and then uses greedy graph coloring to group only compatible tasks into the same optimization step. Every few iterations the scheduler recomputes the graph so the groups adapt as learning proceeds. As our approach can be used in tandem with standard optimizers and add negligible overhead, it can be implemented into existing systems. Experiments on six datasets show consistent gains over strong baselines and stand-alone state-of-the-art approaches.
Our contributions are as follows:
  1. 1.
    We introduce an interference-aware task scheduler based on greedy graph coloring that separates highly conflicting tasks without manual tuning.
  2. 2.
    We build upon and apply existing theory to prove the effectiveness of components within the scheduler
  3. 3.
    We combine the scheduler with PCGrad, AdaTask, and GradNorm and demonstrate systematic improvements across six datasets.
  4. 4.
@@
    We perform ablation studies to prove that dynamic recoloring and history-averaged conflict estimates are essential for high performance.
2 Related Works
Multi-task learning (MTL) methods have evolved from simple loss-weighting approaches to larger and more sophisticated optimization techniques that manage task conflict and cooperation Yang2023AdaTask . Early adaptive-weighting approaches sought to balance losses automatically 9336293 ; Fan2023MaxGNR , while more recent work modifies gradients directly Yu2020PCGrad . Task scheduling and grouping methods, though far less popular than adaptive weighting techniques torbarina2023challengesopportunitiesusingtransformerbased , have contributed to the field by controlling the timing of updates. Our interference-aware scheduler combines these lines of research, offering lightweight, adaptive grouping with provable convergence.
2.1 Tuned Loss Weighting
From early MTL work it became clear that simply summing task losses often favors one objective at the expense of others Kurin2022InDefense ; Zhao2024ExcessMTL ; Mueller2022Temperature , especially when losses have different scales or noise levels. To address this, practitioners manually tuned per‑task weight coefficients ($\lambda$-values) to rebalance learning Liang2020 ; lin2022a , but this process was laborious and dataset-specific. Thus, researchers began to develop automated methods.
2.2 Adaptive Loss Weighting
(Kendall et al., 2018) kendall2018multi introduced uncertainty weighting, learning each task’s homoscedastic (constant‑variance) Bishop2006 noise to scale losses automatically and improve depth and semantics on NYUv2 Silberman2012 .
GradNorm automatically balances multiple loss functions by tuning each task’s gradient magnitude so that all tasks train at comparable speeds Chen2017GradNorm . It does this by introducing a single asymmetry hyperparameter $\alpha$ that governs how much each task’s loss is called. This eliminates the need for expensive grid searchers over manual weights. GradNorm was also a major leap empirically as it surpassed exhaustive search baselines on both regression and classification tasks. Dynamic Weight Averaging (DWA) extended this idea by adjusting weights based on loss rate of change, reducing oscillations between tasks Liu2019EndToEnd .
More recently AdaTask applies task-specific learning rates that adapt to each head’s gradient norm, yielding significant gains on multi-label classification benchmarks Yang2023AdaTask .
2.3 Gradient-Level Conflict Mitigation
Rather than rescaling losses, gradient surgery methods alter update directions. PCGrad projects gradients that conflict (negative cosine) onto each other’s normal plane, significantly boosting efficiency on supervised vision and RL problems Yu2020PCGrad . CAGrad frames task balance as a min-max optimization, finding updates that maximize the worst-case task improvement Liu2021CAGrad . The Multiple Gradient Descent Algorithm (MGDA) computes a Pareto-optimal convex combination of task gradients, ensuring no task is harmed Sener2018 . More recent variants such as SAM-GS incorporate momentum into conflict detection, smoothing gradient estimates while preserving the benefits of surgery Borsani2025SAMGS .
2.4 Empirical Task Grouping
Early approaches under this category used round-robin and random sampling-based approaches that ignored any task relationships McCann2018NaturalLanguageDecathlon ; Zamir2020RobustLearning . (Standley et al., 2020) exhaustively searches over small subsets to identify beneficial groupings, demonstrating the potential of selective updates but failing to scale beyond eight tasks due to computational complexity Standley2020TaskGrouping .
Task Affinity Groupings (TAG) Fifty2021 performs one joint training run to measure inter-task ’affinity’. It quantifies how an update for task $i$ (its gradient) would change task $j$’s loss, and it uses these cross-effects to select partitions of tasks that should share updates. The key idea is to treat grouping as an outcome of measured gradient interactions.
Ayman et al. ayman2023taskgroupingautomatedmultitask train a predictor that maps single-task statistics and dataset features to an estimate of whether two or more tasks should be grouped. They then use that predictor to guide a randomized search over groups, which dramatically reduces the number of multi-task trainings (or ’MTL trials’) needed to find a good partition.
Using a completely different approach, Towards Principled Task Grouping (PTG) wang2024towards formulates grouping as a mathematical program with a theoretically motivated objective capturing beneficial transfer while respecting resources constraints (e.g., compute budgets). It builds a principled optimization over candidate groups that is meant to generalize across application domains.
Scalable Task Grouping via Training Dynamics (STG-MLT) Sherif_2024 avoids expensive affinity estimation by extracting Data Maps swayamdipta-etal-2020-dataset (simple summaries of training dynamics per task) and then clustering tasks using those features. The clusters are intended to push for positive transfer at larger scale. This approach essentially replaces gradient cross-effects with more compact trajectory features that are cheap to compute and easy to cluster.
Selective Task Group Updates (STGU) jeong2025selective moves from static partitions to online grouping with sequential updates. It tracks running estimate of cross-task interaction during training, forms groups, and then updates one group per branch. This couples grouping with the step-to-step training dynamics. The method’s core mechanism is to recompute grouping signals as the model evolves and to partition groups accordingly.
@@ 3 Problem Setup
We formalize multi-task learning (MTL) Caruana1997 as the optimization of a shared network under scheduled task activation. Each task contributes a loss whose gradients may align or conflict. We quantify conflict with the negative-cosine interference coefficient, embed all tasks in a conflict graph, and later use that graph to derive our schedule. This section fixes notation and states the optimization goal that the remainder of the methodology addresses.
@@ 3.1 Data and Notation
Let $\mathcal{T}={T_{1},\dots,T_{K}}$ be the set of learning tasks. Let $\theta\in\mathbb{R}^{d}$ denote shared network parameters and $\phi_{k}\in\mathbb{R}^{d_{k}}$ denote task‑private parameters for $T_{k}$.
Each task provides a dataset $\mathcal{D}_{k}={(x_{i}^{(k)},y_{i}^{(k)})}_{i=1}^{n_{k}}$ and a differentiable loss $\mathcal{L}_{k}(\theta,\phi_{k};\mathcal{D}_{k})$ (e.g., cross-entropy, segmentation losses, etc.). Throughout, we write
$g_{k}=\nabla_{\theta}\mathcal{L}_{k}(\theta,\phi_{k})\quad\text{and}\quad h_{k}=\nabla_{\phi_{k}}\mathcal{L}_{k}(\theta,\phi_{k})$ (1)
for the stochastic gradients obtained from a mini-batch sampled within $\mathcal{D}_{k}$.
@@ 3.1.1 Interference Coefficient
Task interactions are captured by the interference coefficient
$\rho_{ij}\;=\;-\frac{\langle g_{i},g_{j}\rangle}{\|g_{i}\|\,\|g_{j}\|},$ (2)
introduced in PCGrad to detect gradient conflict Yu2020PCGrad and later adopted by CAGrad Liu2021CAGrad and GradNorm variants Chen2017GradNorm . Positive $\rho_{ij}$ indicates that following $g_{i}$ harms $T_{j}$ (negative cosine similarity), whereas $\rho_{ij}\leq 0$ implies alignment or neutrality.
@@ 3.1.2 Conflict Graph
$G_{\tau}=(\mathcal{T},E_{\tau}),\quad E_{\tau}=\bigl\{(i,j)\,:\,\rho_{ij}>\tau\bigr\}.$ (3)
Vertices represent tasks; an edge links any pair whose gradients conflict beyond $\tau$. Greedy graph‑coloring of a graph with maximum degree $\Delta$ uses at most $\Delta+1$ colors bonamy2018boundingchifractiondelta .
@@ 3.2 Goal
At training step $t$ we choose an active task set $S_{t}\subseteq\mathcal{T}$ and perform an SGD update only on those tasks:
[…]
The scheduled MTL objective is therefore
$\min_{\theta,\{\phi_{k}\}}\;F(\theta,\phi)=\sum_{t=1}^{T}\;\sum_{k\in S_{t}}\;\mathcal{L}_{k}\bigl(\theta_{t},\phi_{k,t}\bigr).$ (5)
Our overarching aim is to design the sequence ${S_{t}}_{t=1}^{T}$ so that: (1) Every task is visited regularly (to retain representation quality) and (2) Conflicting tasks seldom appear together (to avoid destructive interference).
[…]
@@ 4 Proposed Approach
We design an interference-aware scheduler that partitions tasks into low-conflict groups and activates exactly one group per optimization step. The procedure consists of four stages: (1) estimating pairwise interference, (2) building and coloring the conflict graph, (3) generating a periodic schedule, and (4) updating that schedule as training evolves.
@@ 4.1 Estimating Gradient Interference
At step $t$ and for every task $T_{k}$ appearing in the current mini‑batch we compute a task‑specific stochastic gradient
@@
using an independent sub‑batch $\mathcal{B}_{k}^{(t)}\subset\mathcal{D}_{k}$. We then update an exponential moving average
@@
Computing all K(K1)/2K(K-1)/2 cosines is $O(K^{2}d)$ with $d$ representing the shared‑parameter dimension.
@@ 4.2 Conflict Graph Construction
$G^{(r)}_{\tau}=(\mathcal{T},E_{\tau}^{(r)}),\quad E_{\tau}^{(r)}=\bigl\{(i,j)\;:\;\rho^{(t_{r})}_{ij}>\tau\bigr\}.$ (9)
Edges connect tasks whose averaged gradients have cosine similarity less than $\tau$. Intuitively, higher $\tau$ yields a sparser graph and longer groups, trading off less conflict for fewer gradient updates per task.
@@ 4.3 Partitioning via Greedy graph coloring
We apply the Welsh-Powell largest-first greedy heuristic Welsh1967 to color $G^{(r)}\tau$ and obtain color classes $C^{(r)}{1},\dots,C^{(r)}_{m_{r}}$. Classical graph-theory results West2000 guarantee the heuristic uses no more than $\Delta+1$ colors, where $\Delta$ is the maximum vertex degree. In practice $\Delta$ is small because many task pairs do not interfere, yielding concise schedules.

### v2

# Graph Coloring for Multi-Task Learning

@@ Abstract
When different objectives conflict with each other in multi-task learning, gradients begin to interfere and slow convergence, thereby potentially reducing the final model’s performance. To address this, we introduce SON-GOKU, a scheduler that computes gradient interference, constructs an interference graph, and then applies greedy graph-coloring to partition tasks into groups that align well with each other. At each training step, only one group (color class) of tasks are activated, and the grouping partition is constantly recomputed as task relationships evolve throughout training. By ensuring that each mini-batch contains only tasks that pull the model in the same direction, our method improves the effectiveness of any underlying multi-task learning optimizer without additional tuning. Since tasks within these groups will update in compatible directions, multi-task learning will improve model performance rather than impede it. Empirical results on six different datasets show that this interference-aware graph-coloring approach consistently outperforms baselines and state-of-the-art multi-task optimizers. We provide extensive theory showing why grouping and sequential updates improve multi-task learning, with guarantees on descent, convergence, and the ability to accurately identify what tasks conflict or align.
@@ 1 Introduction
Multi-task learning (MTL) trains a single model to solve several tasks simultaneously, sharing knowledge across them to learn more effectively (Caruana, 1997). This allows models to generalize better and converge faster. However, a key issue known as negative transfer arises when tasks don’t align very well with each other (Sener & Koltun, 2018; Shi et al., 2023). When two tasks push the shared network in different directions their gradients clash, slowing or even reversing learning. Prior work addresses this issue primarily via (1) gradient manipulation, which reshapes task gradients to reduce conflicts, and (2) loss reweighting, which rescales task objectives to balance their influence. While effective in some specific settings, these strategies typically treat conflict locally at the level of shared-parameter updates and often overlook the evolving global structure of interactions among tasks throughout training.
Some recent works focus on partitioning tasks into subsets (groups) and updating those groups separately. These approaches have been found to improve accuracy and training stability by forming groups with high measured affinity and then updating one group at a time (Fifty et al., 2021; Jeong & Yoon, 2025). Grouping can outperform gradient manipulation and loss reweighting when tasks form clusters with aligned gradients, because each update then reduces direct clashes in the shared layers, lowers gradient variance within the step, and lets compatible tasks reinforce one another while conflicting tasks wait for their turn.
However, grouping methods often face a few key limitations: (1) many rely on dense pairwise affinities that grow noisy and costly as the number of tasks rises (Fifty et al., 2021; Standley et al., 2020; Sherif et al., 2024), and (2) others predetermine or rarely update groups, so they drift as task relations change (Wang et al., 2024; Ruder, 2017), and (3) several use local heuristics that fail to enforce global compatibility or to specify how groups should rotate over time (Zhang & Yang, 2018; Malhotra et al., 2022).
We present SON-GOKU (Scheduling via Optimal INterference-aware Graph-COloring for TasK Grouping in MUltitask Learning). We measure gradient interference, build a graph of tasks from those measurements, greedily color the graph to form non-conflicting compatible task groups, and update one color group per step during training. This design addresses the earlier issues. We estimate the interference graph from lightweight minibatch statistics and keep it sparse, which avoids noisy dense matrices and scales to many tasks. We recolor the graph at regular intervals so the groups track changing relations during training. Greedy graph coloring ensures we update only compatible tasks in each step, and the color order gives a simple way to cycle through the groups. Our proposed scheduler does not have to work in isolation, it can function on top of existing loss-reweighting and gradient-manipulation MTL approaches.
In our theoretical analysis (Section 5) we show that, under standard conditions, SON-GOKU tends to group tasks whose gradients are, on average, aligned within each group, with high probability. We further show that, over a refresh window, sequentially updating these low-conflict groups yields at least as much expected descent as a single mixed update, and strictly more when between-group interference is sufficiently negative. We also prove that SON-GOKU preserves descent and reaches the usual non-convex SGD rate under mild assumptions, with only a small factor that depends on the within-group conflict level. In Appendix D we discuss the scheduler’s amortized time complexity and the tradeoffs it offers between speed and performance. We discuss ways in which practitioners can reduce its time complexity under certain conditions.
Empirical results from experiments demonstrate that SON-GOKU consistently improves outcomes compared to other MTL approaches, especially when SON-GOKU is coupled with existing approaches. Our contributions are as follows:
  • We propose SON-GOKU, an interference-aware scheduler that measures cross-task gradient conflict, builds a conflict graph, colors it to form compatible groups, and activates one group per step. It can be used on top of standard MTL optimizers.
  • We provide theoretical analysis that offers guarantees on SON-GOKU’s grouping, convergence, scheduling behavior, and more.
  • Across six datasets, SON-GOKU improves over strong baselines and pairs well with methods like PCGrad, AdaTask, and GradNorm, delivering consistent gains.
@@
  • We perform an ablation study showing that dynamic recoloring and history-averaged conflict estimates are key contributors to performance.
2 Related Work
Many MTL methods (especially earlier ones) adjust task influence by learning or adapting loss weights. Examples include uncertainty-based scaling (Kendall et al., 2018), rate-based schemes such as DWA (Liu et al., 2019), and fast bilevel formulations like FAMO (Liu et al., 2023). FAMO in particular is notable for its $\mathcal{O}(1)$ per-step time complexity. These approaches keep all tasks active each step while modulating relative magnitudes. A completely different approach, which emerged in 2018 with MGDA (Sener & Koltun, 2018), focuses on updating shared-parameter update directions to mitigate interference. Methods like PCGrad (Yu et al., 2020), CAGrad (Liu et al., 2021), and MGDA (Sener & Koltun, 2018) modify the geometry of the shared update to reduce cross-task conflicts while still updating all tasks each step. A smaller body of work forms subsets of tasks to update together, using offline affinity estimation or training-dynamics signals (Fifty et al., 2021; Standley et al., 2020; Wang et al., 2024; Sherif et al., 2024). Most recently, Selective Task Group Updates proposes online grouping with sequential updates, reporting that update order can influence task-specific learning (Jeong & Yoon, 2025). SON-GOKU differs in mechanism from existing approaches (Section 4). It complements loss reweighting and gradient surgery, and we provide explicit guarantees on descent, convergence, and graph/partition recovery. An expanded discussion of more related work is provided in Appendix M.
@@ 3 Problem Setup
We formalize multi-task learning (MTL) (Caruana, 1997) as optimizing a shared network while activating only a subset of tasks at each step. Each task contributes a loss whose gradients may align or conflict. We quantify conflict using (the negative of) cosine similarity, embed tasks in a conflict graph, and later use that graph to derive a schedule. This section fixes notation and states the optimization goal that the proposed approach addresses.
@@ 3.1 Data and Notation
Let $\mathcal{T}=\{T_{1},\dots,T_{K}\}$ be the set of tasks. The model has shared parameters $\theta\in\mathbb{R}^{d}$ and task-specific parameters $\phi_{k}\in\mathbb{R}^{d_{k}}$ for $T_{k}$. Each task draws examples $(x,y_{k})$ from a distribution $\mathcal{D}_{k}$ and defines a per-example loss $\ell_{k}(\theta,\phi_{k};x,y_{k})$. Its population loss is
$L_{k}(\theta,\phi_{k}):=\mathbb{E}_{(x,y_{k})\sim\mathcal{D}_{k}}\big[\ell_{k}(\theta,\phi_{k};x,y_{k})\big].$ (1)
We minimize the standard weighted MTL objective
$F(\theta,\phi_{1},\dots,\phi_{K})\;=\;\sum_{k=1}^{K}w_{k}\,L_{k}(\theta,\phi_{k}),$ (2)
with nonnegative task weights $w_{k}$ (default $w_{k}=1$). Note that, for simplicity in later sections, we absorb $w_{k}$ into the per-task gradient estimates. This is permissible since positive scalings do not change cosine signs or the induced conflict graph.
At step $t$, for any task $k$ that is active we compute stochastic gradients on a mini-batch $\mathcal{B}_{k}^{(t)}\subset\mathcal{D}_{k}$:
$g_{k}^{(t)}:=\nabla_{\theta}L_{k}(\theta_{t},\phi_{k,t};\mathcal{B}_{k}^{(t)}),\qquad h_{k}^{(t)}:=\nabla_{\phi_{k}}L_{k}(\theta_{t},\phi_{k,t};\mathcal{B}_{k}^{(t)}).$ (3)
In our proposed method, we form exponential moving averages (EMA) of per-task gradients within a refresh window to stabilize cosine estimates so that they do not become stale (Sec. 4).
@@ 3.1.1 Interference Coefficient
We quantify pairwise interaction with the interference coefficient
$\rho_{ij}\;=\;-\frac{\langle\tilde{g}_{i},\tilde{g}_{j}\rangle}{\|\tilde{g}_{i}\|\,\|\tilde{g}_{j}\|},$ (4)
where $\tilde{g}_{i}$ and $\tilde{g}_{j}$ are the EMA-smoothed gradients at refresh. Positive $\rho_{ij}$ indicates conflict (negative cosine). $\rho_{ij}\leq 0$ indicates alignment or neutrality.
@@ 3.1.2 Conflict Graph
$G_{\tau}=(\mathcal{T},E_{\tau}),\qquad E_{\tau}=\bigl\{(i,j):\rho_{ij}>\tau\bigr\}.$ (5)
Vertices are tasks. An edge between a pair means to not update that pair together. We will utilize $G_{\tau}$ for coloring and scheduling in Section 4
@@ 3.2 Goal
At training step $t$ we choose an active set $S_{t}\subseteq\mathcal{T}$ and update only those tasks:
$\theta_{t+1}=\theta_{t}-\eta_{t}\sum_{k\in S_{t}}g_{k}^{(t)},\qquad\phi_{k,t+1}=\begin{cases}\phi_{k,t}-\eta_{t}h_{k}^{(t)},&k\in S_{t},\\ \phi_{k,t},&k\notin S_{t}.\end{cases}$ (6)
The problem the scheduler addresses is to design the sequence $\{S_{t}\}_{t=1}^{T}$ so that: (1) every task is visited regularly; and (2) conflicting tasks seldom appear together. We instantiate this via greedy graph coloring in Section 4 and analyze the guarantees in Section 5.
@@ 4 Proposed Approach
We design an interference-aware scheduler that partitions tasks into low-conflict groups and activates exactly one group per optimization step. The procedure consists of four stages: (1) estimating pairwise interference, (2) building and coloring the conflict graph, (3) generating a periodic schedule, and (4) updating that schedule as training evolves. An overview of the scheduler is provided as Algorithm 1 in Appendix A. A visualization of SON-GOKU is provided in Figure 1 alongside a simple summary in the Figure caption.
@@ 4.1 Estimating Gradient Interference
We absorb task weights into per-task losses, so $g_{k}^{(t)}$ is the gradient of the weighted loss $w_{k}L_{k}$. Cosine calculations and graph construction are not impacted by applying positive scaling.
At step $t$ and for every task $T_{k}$ appearing in the current mini-batch we compute a task‑specific stochastic gradient
@@
using an independent sub-batch $\mathcal{B}_{k}^{(t)}\subset\mathcal{D}_{k}$. We then update an exponential moving average
@@
Computing all K(K1)/2K(K-1)/2 cosines is $O(K^{2}d)$ with $d$ representing the shared‑parameter dimension. We also write $h_{k}^{(t)}=\nabla_{\phi_{k}}\mathcal{L}_{k}\!\bigl(\theta_{t},\phi_{k,t};\mathcal{B}_{k}^{(t)}\bigr)$ for the gradient with respect to the task-specific parameters $\phi_{k}$.
@@ 4.2 Conflict Graph Construction
$G^{(r)}_{\tau}=(V,E_{\tau}^{(r)}),\quad V=\{1,\dots,K\}E_{\tau}^{(r)}=\bigl\{(i,j)\;:\;\rho^{(t_{r})}_{ij}>\tau\bigr\}.$ (10)
To clarify, tasks are indexed by integers $1\ldots K$ in Equation 10. Edges connect tasks whose averaged gradients have cosine similarity less than $\tau$. Intuitively, larger $\tau$ yields a sparser conflict graph, typically fewer colors (larger per-step groups), and more frequent updates per task. Smaller $\tau$ results in a denser graph, more colors (smaller per-step groups), and less frequent updates per task.
@@ 4.3 Partitioning via Greedy Graph Coloring
We apply the Welsh-Powell largest-first greedy heuristic (Welsh & Powell, 1967) to color $G^{(r)}_{\tau}$ and obtain color classes $C^{(r)}_{1},\dots,C^{(r)}_{m_{r}}$. Classical graph-theory results (West, 2000; Diestel, 2017) guarantee the heuristic uses no more than $\Delta+1$ colors, where $\Delta$ is the maximum vertex degree. In practice $\Delta$ is small because many task pairs do not interfere, yielding concise schedules.